OpenStudy (anonymous):

Find three positive consecutive integers such that the product of the first and second is two more than three times the third.

Directrix (directrix):

Let the three consecutive positive integers be the following: x, x + 1, and x + 2 The product of the first two integers is x * (x + 1) Three times the third is 3(x+2) Two more than three times the third is 2 + 3(x+2) Putting these expressions together according to the dictates of the problem: x(x+1) = 2 + 3(x+2) s^2 + x = 2 + 3x + 6 x^2 + x = 8 + 3x x^2 - 2x - 8 = 0 Use the Zero Product Property (x-4)*(x+2) = 0 x = 4 or x = -2 The integers are positive so -2 is excluded. x = 4 x + 1 = 5 x + 2 + 6